The velocity of a body moving along x-axis is given by v =(36+2x)^1/2 where v is in m/s and x is in metre.Find distance travelled by it in 2nd second.
1
Expert's answer
2012-12-26T12:13:41-0500
v=dx/dt
dx/dt=(36+2x)^1/2
dx(36+2x)^-1/2=dt
(1/2)*2*(36+2x)^1/2=t+const
if x(0)=0 36^1/2=const
const=6
x=(t+6)^2/2-18
d=x(2)-x(0)=(64/2-18)-(36/2-18)=14 m